26.414 is 28.1% of what number?
Answer by ewatrrr(24785)

 
Hi
26.414 is 28.1% of what number?
Let x represent that number
28.1% = 28.1/100 = .281
.281*x = 26.414
x = 26.414/.281
x = 94
CHECKING our Answer***
.281*94 = 26.414